OpenStudy (anonymous):

Simplify 10 + 4(x + 1) + 5. a.4x + 6 b.4x + 19 c.4x + 16

Nnesha (nnesha):

1st)distribute parentheses by outside number 2nd) combine like terms

Nnesha (nnesha):

here is an example distribute \[\rm a(b+c)=a \times b+ a \times c =ab+ac\] multiply each term in the parentheses by outside number

Nnesha (nnesha):

so \[4(x+1)= ???\]

OpenStudy (theloshua):